# Measurement as Bricolage: Examining How Data Scientists Construct Target Variables for Predictive Modeling Tasks

### v1

@@ 5.3. Design Implications: Developing a Scaffolding for Effective Bricolage Practice
In addition to re-visiting existing top-down interventions, our findings surface a need for entirely new interventions designed to help data scientists engage in bricolage more effectively. While many toolkits exist to help data scientists engage in inner-loop activities, such as training models (Pang et al., 2020; Dingen et al., 2018; Cabrera et al., 2023a; Yan et al., 2021) and exploring data (Wongsuphasawat et al., 2015; Kery et al., 2018), few support specific bricolage practices highlighted in our findings. Data scientists currently lack tools to help them weigh validity with competing criteria in their formulation, identify and apply strategies, and evaluate validity sub-criteria. We now use our findings as a concrete scaffolding for identifying actionable points of intervention in data scientists’ bricolage practices. These interventions are intended to facilitate specific practices that data scientists found critical during problem formulation, while adding guardrails around others that have the potential to introduce flaws.
@@ 5.3.2. Helping Data Scientists Identify Strategies
Therefore, future HCI and CSCW research might develop systems that help teams exchange knowledge surrounding common (re)formulation challenges and mitigation strategies. While community-based knowledge exchange is central to inner-loop workflows, as illustrated by Stack Overflow (Barua et al., 2014) for coding and Hugging Face (Shen et al., 2024) for model training, data scientists’ support for problem formulation remains siloed. Similar to other online communities, this system might structure discussions via a comment forum where teams share specific strategies—P8’s combining strategy for measuring “academic work ethic” or P6’s bridging strategy for measuring “reading ability.” This forum could eventually be used to inform the design of consensus-backed protocols for navigating common (re)formulation challenges.
@@ 5.3.3. Helping Data Scientists Apply Strategies
Beyond strategy identification, our findings also highlight opportunities to help data scientists apply each strategy more effectively. Therefore, we now describe opportunities to develop tools and processes that support data scientists in applying each strategy while avoiding potential limitations.
@@
Swapping. While swapping, data scientists switch to a different outcome after they identify defects in one or more criteria. Swapping can be adaptive if it helps teams identify a formulation that satisfies all minimum performance standards. Yet swapping can also introduce standard erosion. This tension highlights a need for tools to help data scientists make more systematic swapping decisions. While many existing systems support feature engineering (e.g., (Bhattacharya et al., 2024; Guo et al., 2011; Kandel et al., 2011; Rattenbury et al., 2017)) and EDA (e.g., (Wongsuphasawat et al., 2015; Kery et al., 2018)), only one supports outcome exploration (Gala et al., 2024). HCI and Visualization researchers can support outcome exploration by developing toolkits that help data scientists weigh the portability, simplicity, predictive performance, and resource requirements of alternative outcomes. Because our results suggest that swapping can be enticing given its predictive performance benefits, it is especially important that toolkits help teams identify validity drawbacks of swapping decisions. We discuss opportunities to support validity evaluation in Section 5.3.4.
Bridging. While bridging, data scientists adopt a low-cost, readily available target variable as a proxy for a more costly gold standard measure. When used effectively, bridging is a strong example of bricolage because it involves a creative approach for making do with limited data. Yet the success of bridging hinges on data scientists’ ability to identify trustworthy sources of auxiliary information. To support bridging, data scientists need tools to help them identify, evaluate, and incorporate relevant sources of auxiliary data. While a large body of HCI and Data Engineering research has developed tools supporting data discovery (e.g., (Fernandez et al., 2018; Nargesian et al., 2019; Paton et al., 2023; Bogatu et al., 2020)), existing toolkits are not tailored to support bridging. Future research might close this gap by developing tools that help data scientists forage for relevant auxiliary data and evaluate its expected utility against the time and cost required to incorporate it in their formulation. As with swapping, it is important that such tools help data scientists identify any defects associated with bridging.
@@ 5.3.4. Helping Data Scientists Evaluate Validity
Disciplines in the quantitative social sciences have established numerous guidelines for evaluating the validity of a measurement instrument. Such guidelines are often grouped into validity sub-criteria — such as convergent, divergent, and predictive validity — describing abstract, theoretical assertions about the expected behavior of a measurement instrument (Drost, 2011). Yet our findings illustrate a disconnect between these abstract assertions and the more concrete theory- and data-driven evaluation activities performed by practitioners. We resolve this tension by identifying opportunities to scaffold data scientists’ existing evaluation practices (Table 3).
First, our findings highlight opportunities to help teams incorporate domain knowledge while engaging in theory-driven evaluation activities. For example, HCI, ML, and Visualization researchers might develop toolkits that help teams map hypothesized causal relationships between predictive features, interventions, outcome variables, and latent constructs during problem formulation. In an education context, this tool might help teams chart students’ longitudinal trajectories through the education system, weigh points of intervention, and consider an intervention’s effect on multiple downstream outcomes of interest. This tool might also encourage teams to draw upon their domain knowledge to identify spurious causes of outcome variables — e.g., by listing causes of an outcome that are unrelated to the construct of interest. During the design probe, several participants independently suggested an interest in a tool supporting these theory-driven evaluation activities. While the DoWhy package (Sharma and Kiciman, 2020) enables data scientists to reason about causal relationships in their data, this package is not tailored to support the formulation of prediction tasks.
@@ 5.4. Implications for Generative AI Evaluation
While we study measurement in the context of predictive modeling, our findings also connect to emerging discourse surrounding the evaluation of Generative AI systems. Recent work has cast Generative AI evaluation as a measurement task, in which evaluation designers measure properties of AI systems — such as the “toxicity” or “helpfulness” of outputs — by (1) constructing a systematic definition for a concept to be measured, (2) operationalizing this definition into a measurement instrument (e.g., annotation instructions), and (3) applying this operationalization to obtain measurements (Wallach et al., 2024). Recent empirical studies have also shown that the evaluation design process is deeply iterative: practitioners often require multiple rounds of refinement to reconcile their conceptual understanding of evaluation criteria with measurements obtained on specific model outputs (He et al., 2025; Ashktorab et al., 2024; Szymanski et al., 2024; Shankar et al., 2024; Pan et al., 2024). Teams may balance a range of practical criteria while engaging with this process, such as the extensibility and actionability of measurement instruments (Harvey et al., 2025). More broadly, viewing Generative AI evaluation design as a bricolage practice opens a space of research questions. How do evaluation designers make do when balancing validity with competing criteria under resource constraints (e.g., limited human ratings and real-world usage data)? How might we develop a scaffolding for this iterative process that enhances validity, while also supporting flexibility? Our work offers an entry point to these key questions.
@@ 5.5. Study Limitations
Finally, our findings draw upon data scientists with a baseline level of domain knowledge and technical competency. Our recruitment criteria stipulated that data scientists have at least one year of full-time experience in the healthcare or education sectors to be eligible to participate. Additionally, our quoted examples of (re)formulation strategies in Section 4.2 often feature creative practices by highly experienced practitioners. Less experienced data scientists may draw upon a narrower range of strategies in their repertoire or engage in less sophisticated validity evaluation practices. Understanding and supporting novice practitioners’ bricolage practices is an important area for future research.

### v2

@@ 5.3. Design Implications: Developing a Scaffolding for Effective Bricolage Practice
In addition to re-visiting existing top-down interventions, our findings surface a need for entirely new interventions that center bricolage as a data science practice. While many toolkits help data scientists engage in inner-loop activities, such as training models (Pang et al., 2020; Dingen et al., 2018; Cabrera et al., 2023a; Yan et al., 2021) and exploring data (Wongsuphasawat et al., 2015; Kery et al., 2018), data scientists lack tools to help them weigh validity with competing criteria, identify and apply strategies, and evaluate validity. We now use our findings as a concrete scaffolding for identifying actionable points of intervention in data scientists’ bricolage practices. These interventions are intended to facilitate key practices that data scientists found critical during problem formulation, while adding guardrails around others that have the potential to introduce flaws.
@@ 5.3.2. Helping Data Scientists Identify Strategies
Future HCI and CSCW research might develop systems that help teams exchange knowledge surrounding common (re)formulation challenges. While community-based knowledge exchange is central to inner-loop workflows, as illustrated by Stack Overflow (Barua et al., 2014) for coding and Hugging Face (Shen et al., 2024) for model training, data scientists’ support for problem formulation remains siloed. Similar to other online communities, this system might structure discussions via a comment forum where teams share specific strategies—P8’s combining strategy for measuring “academic work ethic” or P6’s bridging strategy for measuring “reading ability.” This forum could eventually be used to inform the design of consensus-backed protocols for navigating common (re)formulation challenges.
@@ 5.3.3. Helping Data Scientists Apply Strategies
Our findings also highlight opportunities to help data scientists apply each strategy they identify more effectively. Therefore, we now describe opportunities to develop tools and processes that support data scientists in applying each strategy effectively, while also avoiding potential strategy-specific complications.
@@
Swapping. While swapping, data scientists switch to a different outcome after they identify defects in one or more criteria. Swapping can be adaptive if it helps teams identify a formulation that satisfies all minimum performance standards. Yet swapping can also introduce standard erosion. This tension highlights a need for tools to help data scientists make more systematic swapping decisions. While many existing systems support feature engineering (e.g., (Bhattacharya et al., 2024; Guo et al., 2011; Kandel et al., 2011; Rattenbury et al., 2017)) and EDA (e.g., (Wongsuphasawat et al., 2015; Kery et al., 2018)), only one supports outcome exploration (Gala et al., 2024). HCI and Visualization researchers can support outcome exploration by developing toolkits that help data scientists weigh the portability, simplicity, predictive performance, and resource requirements of alternative outcomes. Because our results suggest that swapping can be enticing given its predictive performance benefits, it is especially important that toolkits help teams identify validity drawbacks of swapping decisions. We discuss opportunities to support validity evaluation while considering swapping in Section 5.3.4.
Bridging. While bridging, data scientists adopt a low-cost, readily available target variable as a proxy for a more costly gold standard measure. Bridging exemplifies bricolage because it involves a creative approach for making do with limited data. Yet the success of bridging hinges on data scientists’ ability to identify trustworthy sources of auxiliary information. To support bridging, data scientists need tools to help them identify, evaluate, and incorporate relevant sources of auxiliary data. While HCI and Data Engineering research has developed tools supporting data discovery (e.g., (Fernandez et al., 2018; Nargesian et al., 2019; Paton et al., 2023; Bogatu et al., 2020)), existing toolkits are not tailored to support bridging. Future research might close this gap by developing tools that help data scientists forage for relevant auxiliary data and evaluate its expected utility against the resources required to incorporate it in their formulation. As with swapping, it is important that such tools help data scientists identify any defects associated with bridging.
@@ 5.3.4. Helping Data Scientists Evaluate Validity
Disciplines in the quantitative social sciences have established numerous guidelines for evaluating the validity of a measurement instrument. Such guidelines are often grouped into validity sub-criteria — such as convergent, divergent, and predictive validity — describing abstract, theoretical assertions about the expected behavior of a measurement instrument (Drost, 2011). Yet our findings illustrate a disconnect between these abstract assertions and the more concrete theory- and data-driven evaluation activities performed by practitioners. We resolve this tension by identifying opportunities to more effectively integrate measurement principles into data scientists’ existing evaluation practices (Table 3).
First, our findings highlight opportunities to help teams incorporate domain knowledge while engaging in theory-driven evaluation practices. For example, HCI, ML, and Visualization researchers might develop toolkits that help teams map hypothesized causal relationships between interventions, outcome variables, and latent constructs during problem formulation. In education, this tool might help teams chart students’ longitudinal trajectories through the education system, weigh points of intervention, and consider an intervention’s effect on multiple downstream outcomes. This tool might also encourage teams to draw upon their domain knowledge to identify spurious causes of outcome variables — e.g., by listing causes of an outcome that are unrelated to the construct of interest. During the design probe, several participants suggested an interest in a tool supporting these evaluation practices. While the DoWhy package (Sharma and Kiciman, 2020) enables data scientists to reason about causal relationships in their data, this package is not tailored to support the formulation of prediction tasks.
@@ 5.4. Implications for Generative AI Evaluation
While we study measurement in the context of predictive modeling, our findings also connect to emerging discourse surrounding the evaluation of Generative AI systems. Recent work has cast Generative AI evaluation as a measurement task, in which evaluation designers measure properties of AI systems — such as the “toxicity” or “helpfulness” of outputs — by (1) constructing a systematic definition for a concept, (2) operationalizing this definition into a measurement instrument (e.g., annotation instructions), and (3) applying this operationalization to obtain measurements (Wallach et al., 2024). Recent empirical studies have also shown that the evaluation design process is deeply iterative: practitioners require multiple rounds of refinement to reconcile their conceptual understanding of evaluation criteria with measurements obtained on specific model outputs (He et al., 2025; Ashktorab et al., 2024; Szymanski et al., 2024; Shankar et al., 2024; Pan et al., 2024). Teams may balance a range of practical criteria while engaging with this process, such as the extensibility and actionability of measurement instruments (Harvey et al., 2025). More broadly, viewing Generative AI evaluation design as a bricolage practice opens a space of research questions. How do evaluation designers make do when balancing validity with competing criteria under resource constraints (e.g., limited human ratings and real-world usage data)? How might we develop a scaffolding for this iterative process that enhances validity, while also supporting flexibility? Our work offers an entry point to these key questions.
@@ 5.5. Study Limitations
Finally, our findings draw upon data scientists with a baseline level of domain knowledge and technical competency. Our recruitment criteria stipulated that data scientists have at least one year of full-time experience in the healthcare or education sectors to participate. Additionally, our quoted examples of (re)formulation strategies in Section 4.2 often feature creative practices by highly experienced practitioners. Less experienced data scientists may draw upon a narrower range of strategies in their repertoire or engage in less sophisticated validity evaluation practices. Understanding and supporting novice practitioners’ bricolage practices is an important area for future research.